# see the cls

### v1

@@ Abstract
Complex multi-step reasoning tasks, such as solving mathematical problems or generating code, remain a significant hurdle for even the most advanced large language models (LLMs). Verifying LLM outputs with an Outcome Reward Model (ORM) is a standard inference-time technique aimed at enhancing the reasoning performance of LLMs. However, this still proves insufficient for reasoning tasks with a lengthy or multi-hop reasoning chain, where the intermediate outcomes are neither properly rewarded nor penalized. Process supervision addresses this limitation by assigning intermediate rewards during the reasoning process. To date, the methods used to collect process supervision data have relied on either human annotation or per-step Monte Carlo estimation, both prohibitively expensive to scale, thus hindering the broad application of this technique. In response to this challenge, we propose a novel divide-and-conquer style Monte Carlo Tree Search (MCTS) algorithm named OmegaPRM for the efficient collection of high-quality process supervision data. This algorithm swiftly identifies the first error in the Chain of Thought (CoT) with binary search and balances the positive and negative examples, thereby ensuring both efficiency and quality. As a result, we are able to collect over 1.5 million process supervision annotations to train a Process Reward Model (PRM). Utilizing this fully automated process supervision alongside the weighted self-consistency algorithm, we have enhanced the instruction tuned Gemini Pro model’s math reasoning performance, achieving a 69.4% success rate on the MATH benchmark, a 36% relative improvement from the 51% base model performance. Additionally, the entire process operates without any human intervention, making our method both financially and computationally cost-effective compared to existing methods.
@@ 1 Introduction
Despite significant progress in various large language model (LLM) benchmarks achieved through simply scaling up the model (Wei et al., 2022a), the development of complex reasoning abilities, particularly in tasks like mathematical problem-solving and code generation, necessitates a deeper understanding and remains an active research frontier. Chain-of-Thought (CoT) Prompting (Wei et al., 2022b) was proposed to guide the LLM to break down a reasoning task into a sequence of intermediate steps, similar to a human’s reasoning process. CoT boosts the performance of LLMs on many reasoning tasks, but the greedy decoding strategy limits its performance. To address that, Wang et al. (2023) proposed the self-consistency decoding strategy, leveraging multiple reasoning paths to reach a voted answer. Besides these prompting focused work, fine-tuning LLM with question and CoT solution pairs (Perez et al., 2021; Ouyang et al., 2022) proved to be effective, outperforming prompting-only methods, understandably.
A third active area of research aiming to improve reasoning capability is using a verifier to aid LLM reasoning. Off-the-shelf LLMs are a straightforward option for verifiers (Huang et al., 2022; Luo et al., 2023). While they can be used to verify the result after reaching the final answer, identify potential mistakes and suggest corrections, the performance with such aids is still limited when it comes to complex multi-step math reasoning problems. Since the advent of RLHF (Ouyang et al., 2022), reward models have been used to align LLM behaviors to human preferences and are proven to be a crucial step in training. With reward models as verifiers, model candidate outcomes can be reranked according to the reward to ensure higher accuracy and reliability.
Two predominant types of reward models are frequently employed: the Outcome Reward Models (ORMs; Yu et al., 2024; Cobbe et al., 2021) and the Process Reward Models (PRMs; Li et al., 2023; Uesato et al., 2022; Lightman et al., 2023). The ORM, what’s widely used in RLHF, produces signals only at the end of problem solving, which means the intermediate steps of a reasoning chain are not at all rewarded or penalized. In light of such an obvious shortcoming, the PRM emerged to leverage the reward model in a more granular fashion, by explicitly rewarding or penalizing every reasoning step, namely, with process supervision.
@@
As the name suggests, process supervision requires supervised data. So far, collecting such data relies largely on human annotation (Uesato et al., 2022; Lightman et al., 2023). Annotators have to label each step of every multi-step reasoning problem meticulously, the level of skills and sheer labor demanded render such a practice costly and difficult to scale. Recent efforts to automate such a data collection process with per-step Monte Carlo estimation have shown promising results (Wang et al., 2024a, b). This data collection process excludes the human annotator from the loop, yet the efficiency of such algorithms remains a problem.
Inspired by AlphaGo Zero (Silver et al., 2017), we developed a novel divide-and-conquer style Monte Carlo Tree Search (MCTS) algorithm named OmegaPRM for automated collection of process supervision data. For each question, we build a Monte Carlo Tree, as shown in Fig. 1, with the details explained in Section 3.3. Our algorithm allows for an efficient collection of over 1.5 million high-quality process annotations. Our PRM trained with this dataset, when combined with the weighted self-consistency algorithm to select candidate responses, achieves a 69.4% success rate on the Hendrycks MATH (Hendrycks et al., 2021) benchmark.
@@
  • We combine our verifier with weighted self-consistency to further boost the performance of LLM reasoning. We reached 69.4% success rate on the MATH benchmark.
The rest of the paper is structured as follows. We discuss related work on using LLMs to solve mathematical reasoning problems in Section 2. We describe our main method in Section 3. Our experimental setup including the task, model configuration, baselines and metrics are discussed in Section 4. We showcase our superior result also in Section 4.
@@ 2 Related Work
MCTS (Świechowski et al., 2021) has been widely adopted in reinforcement learning (RL). AlphaGo (Silver et al., 2016) and AlphaGo Zero (Silver et al., 2017) were able to achieve great performance with MCTS and deep reinforcement learning. For LLMs, there are planning algorithms that fall in the category of tree search, such as Tree-of-Thought (Yao et al., 2023), Reasoning-via-Planing (Hao et al., 2023) and inference-time MCTS (Feng et al., 2023). Most of the work focuses on tree-like decoding to find the best output.
@@ 3.1 Process Supervision
Process supervision is a concept proposed to differentiate from outcome supervision. The reward models trained with these objectives are termed Process Reward Models (PRMs) and Outcome Reward Models (ORMs), respectively. In the ORM framework, given a query $q$ (e.g., a mathematical problem) and its corresponding response $x$ (e.g., a model-generated solution), an ORM is trained to predict the correctness of the final answer within the response. Formally, an ORM takes $q$ and $x$ and outputs the probability $p=\mathrm{ORM}(q,x)$ that the final answer in the response is correct. With a training set of question-answer pairs available, an ORM can be trained by sampling outputs from a policy model (e.g., a pretrained or supervised fine-tuned LLM) using the questions and obtaining the correctness labels by comparing these outputs with the golden answers.
@@ 3.2 Process Annotation with Monte Carlo Method
(a) Monte Carlo estimation of a prefix solution.
(b) Error locating using binary search.
(c) Three stages in an iteration of the MCTS process.
@@ 3.4 PRM Training
where $\hat{y}_{i}$ represents the correctness label and $y_{i}=\mathrm{PRM}(s,a)$ is the prediction score of the PRM. Wang et al. (2024b) have used the Monte Carlo estimation as the correctness label, denoted as $\hat{y}=\mathrm{MC}(s)$. Alternatively, Wang et al. (2024a) have employed a binary labeling approach, where $\hat{y}=\mathbf{1}[\mathrm{MC}(s)>0]$, assigning $\hat{y}=1$ for any positive Monte Carlo estimation and $\hat{y}=0$ otherwise. We refer the former option as pointwise soft label and the latter as pointwise hard label. In addition, considering there are many cases where a common solution prefix has multiple single-step actions, we can also minimize the cross-entropy loss between the PRM predictions and the normalized pairwise preferences following the Bradley-Terry model (Christiano et al., 2017). We refer this training method as pairwise approach. We use the pointwise soft label when evaluating the main results in Section 4.1, and a comparion of the three objectives are discussed in Section 4.3.
@@ 4 Experiments
We conduct our experiments on the challenging MATH dataset (Hendrycks et al., 2021). We use the same training and testing split as described in Lightman et al. (2023), which consists of $12$K training examples and a subset with $500$ representative problems from the original $4.5$K testing examples introduced in Hendrycks et al. (2021). We observe similar policy performance on the full test set and the subset. For creating the process annotation data, we use the questions from the training split and set the search limit to $100$ per question, resulting $1.5$M per-step process supervision annotations. We use $\alpha=0.5$, $\beta=0.9$ and $L=500$ for calculating $Q(s,r)$ in Eq. 2; and $c_{\text{puct}}=0.125$ in Eq. 3.
Base model.
As mentioned in previous studies (Lightman et al., 2023; Wang et al., 2024a, b), the quality of the policy model significantly influences the quality of sampled model responses, thereby affecting the PRM data quality. Lightman et al. (2023) obtain their base model by fine-tuning GPT-4 (OpenAI, 2023) on MathMix, a corpus of $1.5$B math-related tokens. Meanwhile, Wang et al. (2024a) employed various base models fine-tuned on MetaMath (Yu et al., 2023) for their policy models. Similarly, we tune a pretrained Gemini Pro (Gemini Team et al., 2023) by distilling knowledge from Gemini Ultra on math instruction datasets, resulting in a policy model that achieves an accuracy of approximately $51$% on the MATH test set. This results in a policy model significantly smaller than GPT-4 used in Lightman et al. (2023), enabling more efficient data sampling while maintaining competitive math problem-solving capabilities. All the PRMs are trained based on the pretrained Gemini Pro.
@@
We evaluate the PRM-based majority voting results following Wang et al. (2024a) using PRMs trained on different datasets. The final solution score is calculated by the product of scores across all steps. Baseline datasets include PRM800K (Lightman et al., 2023) and Math-Shepherd (Wang et al., 2024a), both publicly available. Additionally, we generate a process annotation dataset with our Gemini policy model using the brute-force approach described in Wang et al. (2024b), referred to as MiPS in subsequent sections.
@@ 4.1 Main Results
Figure 3: A comparison of PRMs trained with different process supervision datasets, evaluated by their ability to search over many test solutions using a PRM-weighted majority voting. We visualize the variance across many sub-samples of the $256$ solutions we generated in total per problem.
Fig. 3 presents the performance comparison of PRMs trained on various process annotation datasets. OmegaPRM consistently outperforms the other datasets. Specifically, the fine-tuned Gemini Pro achieves $69.4$% accuracy on MATH using OmegaPRM-weighted majority voting. It shows superior performance comparing to both human annotated PRM800K but also automatic annotated Math-Shepherd and MiPS. More specifically, when the number of samples is small, all the PRM models outperformed the majority vote. However, as the number of samples increases, the performance of other PRMs gradually converges to the same level of the majority vote. In contrast, our PRM model continues to demonstrate a clear margin of accuracy.
@@ 5 Limitations
Our method for automatic process supervision annotation introduces some noise because false positives and negatives, but experiments indicate that it can still effectively train a PRM. The PRM trained on our dataset performs better than one trained on the human-annotated PRM800K dataset. The precise impact of noise on PRM performance remains uncertain. For future research, a comprehensive comparison of human and automated annotations should be conducted. One other idea is to integrate human and automated annotations, which could result in more robust and efficient process supervision.

### v2

@@ Abstract
Complex multi-step reasoning tasks, such as solving mathematical problems or generating code, remain a significant hurdle for even the most advanced large language models (LLMs). Verifying LLM outputs with an Outcome Reward Model (ORM) is a standard inference-time technique aimed at enhancing the reasoning performance of LLMs. However, this still proves insufficient for reasoning tasks with a lengthy or multi-hop reasoning chain, where the intermediate outcomes are neither properly rewarded nor penalized. Process supervision addresses this limitation by assigning intermediate rewards during the reasoning process. To date, the methods used to collect process supervision data have relied on either human annotation or per-step Monte Carlo estimation, both prohibitively expensive to scale, thus hindering the broad application of this technique. In response to this challenge, we propose a novel divide-and-conquer style Monte Carlo Tree Search (MCTS) algorithm named OmegaPRM for the efficient collection of high-quality process supervision data. This algorithm swiftly identifies the first error in the Chain of Thought (CoT) with binary search and balances the positive and negative examples, thereby ensuring both efficiency and quality. As a result, we are able to collect over 1.5 million process supervision annotations to train Process Reward Models (PRMs). This fully automated process supervision alongside the weighted self-consistency algorithm is able to enhance LLMs’ math reasoning performances. We improved the success rates of the instruction-tuned Gemini Pro model from 51% to 69.4% on MATH500 and from 86.4% to 93.6% on GSM8K. Similarly, we boosted the success rates of Gemma2 27B from 42.3% to 58.2% on MATH500 and from 74.0% to 92.2% on GSM8K. The entire process operates without any human intervention or supervision, making our method both financially and computationally cost-effective compared to existing methods.
@@ 1 Introduction
Despite the impressive advancements achieved by scaling Large Language Models (LLMs) on established benchmarks (Wei et al., 2022a), cultivating more sophisticated reasoning capabilities, particularly in domains like mathematical problem-solving and code generation, remains an active research area. Chain-of-thought (CoT) generation is crucial for these reasoning tasks, as it decomposes complex problems into intermediate steps, mirroring human reasoning processes. Prompting LLMs with CoT examples (Wei et al., 2022b) and fine-tuning them on question-CoT solution pairs (Perez et al., 2021; Ouyang et al., 2022) have proven effective, with the latter demonstrating superior performance. Furthermore, the advent of Reinforcement Learning with Human Feedback (RLHF; Ouyang et al., 2022) has enabled the alignment of LLM behaviors with human preferences through reward models, significantly enhancing model capabilities.
@@
Beyond prompting and further training, developing effective decoding strategies is another crucial avenue for improvement. Self-consistency decoding (Wang et al., 2023) leverages multiple reasoning paths to arrive at a voted answer. Incorporating a verifier, such as an off-the-shelf LLM (Huang et al., 2022; Luo et al., 2023), can further guide LLMs in reasoning tasks by providing a feedback loop to verify final answers, identify errors, and suggest corrections. However, the gain of such approaches remains limited for complex multi-step reasoning problems. Reward models offer a promising alternative to verifiers, enabling the reranking of candidate outcomes based on reward signals to ensure higher accuracy. Two primary types of reward models have emerged: Outcome Reward Models (ORMs; Yu et al., 2024; Cobbe et al., 2021), which provide feedback only at the end of the problem-solving process, and Process Reward Models (PRMs; Li et al., 2023; Uesato et al., 2022; Lightman et al., 2023), which offer granular feedback at each reasoning step. PRMs have demonstrated superior effectiveness for complex reasoning tasks by providing such fine-grained supervision.
The primary bottleneck in developing PRMs lies in obtaining process supervision signals, which require supervised labels for each reasoning step. Current approaches rely heavily on costly and labor-intensive human annotation (Uesato et al., 2022; Lightman et al., 2023). Automating this process is crucial for scalability and efficiency. While recent efforts using per-step Monte Carlo estimation have shown promise (Wang et al., 2024a, b), their efficiency remains limited due to the vast search space. To address this challenge, we introduce OmegaPRM, a novel divide-and-conquer Monte Carlo Tree Search (MCTS) algorithm inspired by AlphaGo Zero (Silver et al., 2017) for automated process supervision data collection. For each question, we build a Monte Carlo Tree, as shown in Fig. 1, with the details explained in Section 3.3. This algorithm enables efficient collection of over 1.5 million high-quality process annotations without human intervention. Our PRM, trained on this dataset and combined with weighted self-consistency decoding, significantly improves the performance of instruction-tuned Gemini Pro from 51% to 69.4% on MATH500 (Lightman et al., 2023) and from 86.4% to 93.6% on GSM8K (Cobbe et al., 2021). We also boosted the success rates of Gemma2 27B from 42.3% to 58.2% on MATH500 and from 74.0% to 92.2% on GSM8K.
@@
  • We combine our verifier with weighted self-consistency to further boost the performance of LLM reasoning. We significantly improves the success rates from 51% to 69.4% on MATH500 and from 86.4% to 93.6% on GSM8K for instruction-tuned Gemini Pro. For Gemma2 27B, we also improved the success rates of from 42.3% to 58.2% on MATH500 and from 74.0% to 92.2% on GSM8K.
@@ 2 Related Work
MCTS (Świechowski et al., 2021) has been widely adopted in reinforcement learning (RL). AlphaGo (Silver et al., 2016) and AlphaGo Zero (Silver et al., 2017) were able to achieve great performance with MCTS and deep reinforcement learning. For LLMs, there are planning algorithms that fall in the category of tree search, such as Tree-of-Thought (Yao et al., 2023) and Reasoning-via-Planing (Hao et al., 2023). Recently, utilizing tree-like decoding to find the best output during the inference-time has become a hot topic to explore as well, multiple works (Feng et al., 2023; Ma et al., 2023; Zhang et al., 2024; Tian et al., 2024; Feng et al., 2024; Kang et al., 2024) have observed improvements in reasoning tasks.
@@ 3.1 Process Supervision
Process supervision is a concept proposed to differentiate from outcome supervision. The reward models trained with these objectives are termed Process Reward Models (PRMs) and Outcome Reward Models (ORMs), respectively. In the ORM framework, given a query $q$ (e.g., a mathematical problem) and its corresponding response $x$ (e.g., a model-generated solution), an ORM is trained to predict the correctness of the final answer within the response. Formally, an ORM takes $q$ and $x$ and outputs the probability $p=\mathrm{ORM}(q,x)$ that the final answer in the response is correct. With a training set of question-answer pairs available, an ORM can be trained by sampling outputs from a policy model (e.g., a pretrained or fine-tuned LLM) using the questions and obtaining the correctness labels by comparing these outputs with the golden answers.
@@ 3.3 Monte Carlo Tree Search
(a) Monte Carlo estimation of a prefix solution.
(b) Error locating using binary search.
(c) Three stages in an iteration of the MCTS process.
@@ 3.4 PRM Training
where $\hat{y}_{i}$ represents the correctness label and $y_{i}=\mathrm{PRM}(s,a)$ is the prediction score of the PRM. Wang et al. (2024b) have used the Monte Carlo estimation as the correctness label, denoted as $\hat{y}=\mathrm{MC}(s)$. Alternatively, Wang et al. (2024a) have employed a binary labeling approach, where $\hat{y}=\mathbf{1}[\mathrm{MC}(s)>0]$, assigning $\hat{y}=1$ for any positive Monte Carlo estimation and $\hat{y}=0$ otherwise. We refer the former option as pointwise soft label and the latter as pointwise hard label. In addition, considering there are many cases where a common solution prefix has multiple single-step actions, we can also minimize the cross-entropy loss between the PRM predictions and the normalized pairwise preferences following the Bradley-Terry model (Christiano et al., 2017). We refer this training method as pairwise approach, and the detailed pairwise loss formula can be found in Section Appendix B.
We use the pointwise soft label when evaluating the main results in Section 4.1, and a comparion of the three objectives are discussed in Section 4.3.
@@ 4 Experiments
We conduct our experiments on the challenging MATH dataset (Hendrycks et al., 2021). We use the same training and testing split as described in Lightman et al. (2023), which consists of $12$K training examples and a subset with $500$ holdout representative problems from the original $5$K testing examples introduced in Hendrycks et al. (2021). We observe similar policy performance on the full test set and the subset. For creating the process annotation data, we use the questions from the training split and set the search limit to $100$ per question, resulting $1.5$M per-step process supervision annotations. To reduce the false positive and false negative noise, we filtered out questions that are either too hard or too easy for the model. Please refer to Appendix A for details. We use $\alpha=0.5$, $\beta=0.9$ and $L=500$ for calculating $Q(s,r)$ in Eq. 2; and $c_{\text{puct}}=0.125$ in Eq. 3. We sample $k=8$ rollouts for each Monte Carlo estimation.
Models.
In previous studies (Lightman et al., 2023; Wang et al., 2024a, b), both proprietary models such as GPT-4 (OpenAI, 2023) and open-source models such as Llama2 (Touvron et al., 2023) were explored. In our study, we perform experiments with both proprietary Gemini Pro (Gemini Team et al., 2024) and open-source Gemma2 (Gemma Team et al., 2024) models. For Gemini Pro, we follow Lightman et al. (2023); Wang et al. (2024a) to initially fine-tune it on math instruction data, achieving an accuracy of approximately $51$% on the MATH test set. The instruction-tuned model is then used for solution sampling. For open-source models, to maximize reproducibility, we directly use the pretrained Gemma2 27B checkpoint with the 4-shot prompt introduced in Gemini Team et al. (2024). The reward models are all trained from the pretrained checkpoints.
@@
We evaluate the PRM-based majority voting results on GSM8K (Cobbe et al., 2021) and MATH500 (Lightman et al., 2023) using PRMs trained on different process supervision data. We choose the product of scores across all steps as the final solution score following Lightman et al. (2023), where the performance difference between product and minimum of scores was compared and the study showed the difference is minor. Baseline process supervision data include PRM800K (Lightman et al., 2023) and Math-Shepherd (Wang et al., 2024a), both publicly available. Additionally, we generate a process annotation dataset with our Gemini policy model using the brute-force approach described in Wang et al. (2024a, b), referred to as Math-Shepherd (our impl) in subsequent sections.
@@ 4.1 Main Results
(a) Gemini Pro on MATH500.
(b) Gemini Pro on GSM8K.
(c) Gemma2 27B on MATH500.
(d) Gemma2 27B on GSM8K.
Figure 3: A comparison of PRMs trained with different process supervision datasets, evaluated by their ability to search over many test solutions using a PRM-weighted majority voting. We visualize the variance across many sub-samples of the $128$ solutions we generated in total per problem.
Table 1: The performance comparison of PRMs trained with different process supervision datasets. The numbers represent the percentage of problems solved using PRM-weighted majority voting with $k=64$.
MATH500 GSM8K
Gemini Pro Gemma 2 27B Gemini Pro Gemma 2 27B
MajorityVote@64 67.2 54.7 92.7 90.6
   + Math-Shepherd 67.2 57.4 92.7 90.5
   + Math-Shepherd (our impl) 67.2 55.2 91.8 91.4
   + PRM800K 67.6 57.2 92.9 91.7
   + OmegaPRM 69.4 58.2 93.6 92.2
Table 1 and Fig. 3 presents the performance comparison of PRMs trained on various process annotation datasets. OmegaPRM consistently outperforms the other process supervision datasets. Specifically, the fine-tuned Gemini Pro achieves $69.4$% and $93.6$% accuracy on MATH500 and GSM8K, respectively, using OmegaPRM-weighted majority voting. For the pretrained Gemma2 27B, it also performs the best with $58.2$% and $92.2$% accuracy on MATH500 and GSM8K, respectively. It shows superior performance comparing to both human annotated PRM800K but also automatic annotated Math-Shepherd. More specifically, when the number of samples is small, almost all the PRM models outperforme the majority vote. However, as the number of samples increases, the performance of other PRMs gradually converges to the same level of the majority vote. In contrast, our PRM model continues to demonstrate a clear margin of accuracy.
@@ 4.3 PRM Training Objectives
4.4 Algorithm Efficiency
As described in Section Section 3.2 and Section 3.3, we utilize binary search and Monte Carlo Tree Search to improve the efficiency of OmegaPRM process supervision data collection by effectively identifying the first incorrect step and reusing rollouts in Monte Carlo estimation. To quantitatively measure the efficiency of OmegaPRM, we collected process supervision data using both brute-force-style method (Wang et al., 2024a, b) and OmegaPRM with the same computational budget. As a result, we were able to generate $200$K data points using the brute-force algorithm compared to $15$ million data points with OmegaPRM, demonstrating a $75$-times efficiency improvement. In practice, we randomly down-sampled OmegaPRM data to $1.5$ million for PRM training.
@@ 5 Limitations
Our method for automatic process supervision annotation introduces noise in the form of false positives and negatives, but experiments indicate that it can still effectively train a PRM. The PRM trained on our dataset performs better than one trained on the human-annotated PRM800K dataset. The precise impact of noise on PRM performance remains uncertain. For future research, a comprehensive comparison of human and automated annotations should be conducted. One other idea is to integrate human and automated annotations, which could result in more robust and efficient process supervision.
@@ 7 Acknowledgements
Appendix
Appendix A Question Filtering
During the evaluation of partial solution correctness using MC estimation, false negative noise may be introduced when a question is too hard for the model, thus no correct rollout can be found even with correct partial solution. Or false positive noise may be introduced when a question is too easy, that model can conclude in correct answer given partial solution with wrong step. It is not possible to exclude such noise completely, but we can reduce the chance by filtering out questions that are either too hard or too easy for the model. Specifically, we ran a $k=32$ rollouts for each question in the 12K training data, and filter out the questions that with no correct answer (too hard) or no wrong answer (too easy) in the 32 rollouts.
Appendix B Pairwise Loss Formula
When training with pairwise labels, the Bradley-Terry model (people typically use this objective to train reward models in RLHF) generally accepts two probability scalars summing up to 1. When we select the two actions as a pair, there are two cases. The first case is that one sample with a zero MC value, and the other sample with a positive MC value. The second case is that both samples are with positive MC values. The first case is straight-forward, and a normalization step is required for the second case.
Assume the two MC values are $p$ and $q$, and they follow the Bernoulli distribution: $P(X=1)=p$ and $P(Y=1)=q$. We need to calculate the probability that action X is preferred over action Y and vice versa.
$\begin{split}P(X>Y)&=P(X=1,Y=0)=p(1-q),\\ P(X<Y)&=P(X=0,Y=1)=(1-p)q,\\ P(X=Y)&=P(X=0,Y=0)+P(X=1,Y=1)=(1-p)(1-q)+pq.\end{split}$ (5)
For the tied situation, each action has half the chance of being preferred. Thus,
$\begin{split}P(\text{action X is preferred})=P(X>Y)+1/2*P(X=Y)=1/2*(1+p-q),\\ P(\text{action Y is preferred})=P(X<Y)+1/2*P(X=Y)=1/2*(1+q-p).\end{split}$ (6)
Now the MC values are normalized and we can train with the pairwise loss.